# Progressive Checkerboards for Autoregressive Multiscale Image Generation

### v1

@@ 1 Introduction
Image generation methods, including autoregressive (AR) models, diffusion and masked models [24, 23, 29, 38, 32, 25, 9, 27, 2, 42], have shown impressive results in generating high-quality images. Recently, scalewise autoregressive models [37, 5, 18] condition from coarse to fine as their progressive sequence, while parallel AR models [40, 43] select multiple locations at once and condition using sequential blocks.
Both conditioning methods aim to model dependencies between sampling locations. Independently sampling mutually dependent locations can mix modes and produce incompatible values. This is particularly true for adjacent and nearby locations: if an image patch is red, the one next to it is likely to be red as well. Conditioning each location on a common parent, for example in the previous scale, makes the conditional samples more independent of each other [37].
Relying exclusively on scalewise conditioning, however, critically depends on a slow scale-up factor: If the scale factor is too large, then an object spanning multiple locations in one scale may not yet be visible in the previous scale, so its appearance dependencies would not be conditioned and independently sampling can mix modes. Rather than scaling up by a factor of 2, as one might consider initially, the best performing multiscale models [37, 5, 18] currently scale by a factor of $\sqrt[3]{2}\approx 1.26$ (i.e. each factor of 2 is further subdivided into 3 evenly spaced scales). However, slow scale-up is not the only way to insert the necessary conditioning. As prior sequential and parallel methods [40, 43] have shown, another way is to condition between sampling locations.
In this paper, we examine conditioning between locations within each scale, making use of both between- and within-scale conditioning. We develop a sampling order based on a progressive checkerboard, which reduces mutual dependence within each sampling block and enables fast scale-up. Our ordering maintains balance at all levels in a quadtree subdivision, so that varying the block size effectively varies the trade-off between parallelism and modeling conditional dependencies. We use this to explore the relationships between scale-wise and within-scale conditioning, finding somewhat surprisingly that for this spatially balanced setup, the total number of sequential steps matters much more than the scaling factor. Our model’s performance is competitive with state-of-the-art methods in ImageNet 256x256 class-conditioned generation, using fewer sampling steps.
@@ 2.1 Scalewise Autoregression
Gradual scaling AR models [37, 18, 5], first introduced by Tian et al. [37], progressively develop a full-resolution array of latent codes, which they use as a working “canvas” to apply residuals from each scale. At each iteration, the canvas is downsampled to the current scale (which grows by a factor of $\sqrt[3]{2}$ each iteration) and a transformer predicts residuals, which are upsampled and applied back to the full-resolution array.
@@
Viewing the method this way sheds some light on reasons why the scaling factor remains small. Just as diffusion models struggle with large denoising steps due to their use of independent sampling and Gaussian noise [33, 31, 34, 17], scalewise AR models can struggle with large deblurring steps due to independently sampling underlying multi-modal distributions [1, 11].
@@ 2.2 Parallel Sampling
Parallel autoregression (PAR) by Wang et al. [40] generates images by sampling in parallel from equal-sized square partitions. Their method uses raster order within each partition and no multi-scale sampling, limiting its parallelism to only four groups. Our multi-scale checkerboard sampling uses both multi-scale and within-scale conditioning, as well as an efficient ordering that maintains sample balance and spatial diversity at all levels, reducing the number of sequential steps to just 17, while realizing further performance gains.
More recently, Zhang et al. [43] use a locality-aware ordering that adds far-away locations while simultaneously growing already-selected regions with adjacent samples, while Pang et al. [19] use random orderings. However, the former uses a more complex dynamic evolution to grow out regions, while both methods require additional position probe tokens, extending the overall sequence length. We use a simple but effective regular pattern combined with multiscale conditioning, and do not require additional probe tokens.
Yan et al. [41] use a two-stage "generation then reconstruction" approach applied in the context of MAR [15] that restricts the unmasking schedule to far-away locations using striations, the final stage of which corresponds to a mod 2 checkerboard. In contrast to their method, we use a fully balanced progressive checkerboard for all steps, along with explicit multiscale conditioning, in the context of direct autoregression instead of masking. Additional recent works include NAR [6], which uses different decoder heads for horizontal and vertical conditioning, leading to a striation pattern, and ARPG [14], which uses cross-attention for conditioning in random-ordered parallel sampling with masked models.
Some gradual scaling models also incorporate ways to reduce same-scale dependencies, though to a lesser extent. Han et al. [5] use random bit flips (“bitwise self-correction”) to address errors from independent sampling at the quantizer bit level after incorporating BSQ [44] as their tokenizer. Ma et al. [18] use a miniature masked image model applied in the autoregressor head to condition lower-confidence locations on higher-confidence ones. Somewhat further afield, to enable 2x scaling factors, Ren et al. [26] stack a flow-matching diffusion model on top of mixed-mode AR samples to resolve interdependences. Kutscher et al. [12] examine the impact of patch ordering in the context of recognition models.
@@ 3.1 Autoregressor
Figure 2: Overview of our multiscale blockwise checkerboard autoregressor. (For illustration there are $P=4$ checkerboard blocks, while our actual model uses $P=8$ for more spatial distancing.)
Our multiscale progressive checkerboard autoregressive model is built around a transformer [39, 21, 22] with blockwise causal mask using successive checkerboard sampling blocks.
@@
During training, the entire sequence is processed in parallel using the ground truth codes $z_{s}$ for both inputs and targets in a parallel autoregressive fashion. We use a blockwise causal mask, where all tokens within each block $b_{s}^{i}$ can attend to each other, as well as to tokens from all previous blocks, both for the same scale and in lower-resolution scales. Each token position outputs softmax logits for a quantized codebook; we use flat mean cross-entropy loss over all scales and positions.
Importantly, the locations in each block are spaced as evenly as possible. The number of locations that can be sampled in parallel depends on the independence between locations, when conditioned both on the previous scale and previous samples. Thus, there is a trade-off between the number of scales and the degree of parallelism within each scale, a relation that we explore later in Sec 4.3.
@@ 3.2 Progressive Scan Order
To create the progressive checkerboard ordering, we use a balanced quadtree index generator, where the 2D grid of size $N\times N$ is recursively subdivided into quadrants and indices are assigned in such a way as to maintain balanced assignment at all quadtree levels.
We use a divide-and-conquer approach, where at each recursion level, balanced index lists are first generated for each quadrant recursively, then merged using round-robin selection with a diagonal skip-step pattern (i.e., TL, BR, TR, BL). This results in a spatially balanced progressive order, with a unique index in $1\dots N^{2}$ assigned to each of the $N\times N$ spatial positions. See Algorithm 1.
This scan order can be used flexibly with various partition sizes. To create the blocks $b_{s}^{i}$ for each scale $s$, we simply divide the progressive scan order into $P$ contiguous segments of equal size. If the grid size $N$ is not a power of 2 (as is the case in our scaling experiments), we generate the order on the next power of 2 and restrict to the grid size. Figure 3 shows an example progressive checkerboard ordering on an $8\times 8$ grid.
@@ 3.3 Token Embeddings
For token embeddings, we use quantized latent codes from a VAE-based autoencoder [24, 27, 13, 4]. Each scale is encoded separately from the RGB image, so that each scale’s latents are directly interpretable and decodable. For transformer inputs, we pass the VAE latent codes through a small MLP to unfold them from their low-dimensional clustering space back to higher dimension. For outputs, we also tried using the cluster vectors in an MLP mapping to the weights of the linear output head; however, we found no benefit compared to ordinary linear head based on the code indices.
As described in section 3.1, for inputs, we linearly combine the upsampled latents from the previous scale with the shifted outputs from the current scale. This allows the transformer to condition on both sources of information in a pipelined fashion. In addition, we use a class token as the first input, which along with AdaLN [20] is used to condition on class labels with classifier-free guidance [10].
@@ 3.4 Position Encodings and RoPE Mixing
To make use of our blockwise checkerboard order’s ability to model local dependencies via sequential sampling, the transformer must be able to identify and attend to adjacent locations for each token. We accomplish this using both input position embeddings and learned rotary position encodings (RoPE) [35, 7] in self-attention layers.
We initialize RoPE encodings using factored space and scale frequency representations (7/8 spatial, 1/8 scale). Note the embeddings encode the position in the image-scale space, not the sampling order. The embeddings are then learned in the joint space, allowing the model to find a high-frequency adjacent-position basis and other attention patterns optimized for our sampling order.
Note each token in the transformer processes inputs from two locations: one for the current sampling positions $b_{s}^{i}$, and one for conditioning latents at the previous block’s positions $b_{s}^{i-1}$ (Sec. 3.1). Eq. (1) supplies both in the linear mixture for the input layer. However, without modification, the attention layers only use RoPEs for the current positions $b_{s}^{i}$. To incorporate both sets of positions, we looked at mixing RoPEs for the attention keys using learned mixing coefficients [^1] We learn the “logits” $\alpha^{\prime}$ of the mixing coefficients and set $\alpha_{lh}={\rm sigmoid}(\alpha^{\prime}_{lh})$. $\alpha_{lh}$ for each transformer layer $l$ and attention head $h$:
@@ 4.1 Model and Training Details
We train our models on ImageNet [28] at 256x256 resolution. We experiment with two model sizes: the small (S) model has 12 transformer layers with hidden dimension 512 and 16 attention heads, while the large (L) model has 20 layers with hidden dimension 1024 and 16 heads. In most ablation experiments we use the S model size, trained for 100 epochs with a batch size of 128 and learning rate $1\times 10^{-4}$, then dropped to $1\times 10^{-5}$ in two steps over 10 epochs. The L model is trained for 200 epochs with batch size 64 and learning rate $5\times 10^{-5}$, then dropped to $1\times 10^{-5}$ for 5 epochs, followed by a second effective drop by increasing batch size to 320 via gradient accumulation for the last 5 epochs. In all cases, we use AdamW [16] optimizer with weight decay 0.01 and ten-crop transforms, running on a single NVIDIA GH200 GPU.
We use the VAE autoencoder from LlamaGen [36], with a codebook size of 4096. To better represent the smaller image sizes in our multiscale setting, we fine-tune by freezing all layers other than the quantizer codebook itself, and retrain the codebook layer on ImageNet for one epoch with random image sizes between 16 and 256, using an L2 loss on the latent codes before and after quantization.
To measure sample quality, we use Frechet Inception Distance (FID) [8] and Inception Score (IS) [30], computed using 50k samples against the full ImageNet validation set using the TensorFlow implementation from [36, 3].
For classifier-free guidance (CFG), we found applying CFG too early in the sampling progression limits sample diversity, likely due to the first samples’ influence on global structure. Because of this, we sample the first 5 steps (corresponding to the 1x1 and 2x2 scales when scaling by 2x) with CFG=0, and apply CFG for all subsequent steps. Following common practice [40, 6, 43], we perform a sweep of CFG values at 0.1 increments.
@@ 4.2 Benchmark Comparison
LlamaGen-L [36] 343M 3.07 256.1 0.83 0.52 576 12.58
VAR-d16 [37] 310M 3.30 274.4 0.84 0.51 10 0.12
PAR-L-4x [40] 343M 3.76 218.9 0.84 0.50 147 3.38
RandAR-L [19] 343M 2.55 288.8 0.81 0.58 88 1.97
NAR-L [6] 372M 3.06 263.9 0.81 0.53 31 1.01
ARPG-L [14] 320M 2.30 297.7 0.82 0.56 32 0.58
LPD-L [43] 337M 2.40 284.5 0.81 0.57 20 0.28
Checkerboard-L 2x cfg=1.4 (Ours) 343M 2.72 302.5 0.81 0.56 17 0.52
Checkerboard-L 2x cfg=1.5 (Ours) 343M 2.83 318.2 0.82 0.57 17 0.52
Checkerboard-L 4x cfg=1.7 (Ours) 343M 2.79 311.5 0.80 0.57 17 0.52
Table 1: Autoregressive image generation models on ImageNet 256x256. Inference time measured for single image generation on A100.
@@
Table 1 compares our method against recent autoregression-based image generation models of similar model size on ImageNet 256x256. We show results for our Checkerboard-L model with a 2x scaling ratio and 4 sampling steps per scale (17 total), as well as a 4x ratio with 8 steps per scale (also 17 total) [^2] For the L models, the 2x model is trained from scratch for 200 epochs, and the 4x model is initialized with the 2x model weights and trained for 30 epochs.. Compared to PAR [40] and RandAR [19], to which our method is most closely related, we achieve similar or better FID and IS with fewer sampling steps and faster inference time. While PAR uses 147 steps and RandAR 88 steps, we only need 17 steps to achieve a FID of 2.72 (compared to 3.76 and 2.55, respectively). Our inference time of 0.52s per image is also faster than both methods (3.38s for PAR and 1.97s for RandAR) on A100. In addition, our method is competitive with the more recent AR-based models ARPG [14] and LPD [43], notably also using fewer steps.

### v2

@@ 1 Introduction
Image generation methods, including autoregressive (AR) models, diffusion and masked models, have shown impressive results in generating high-quality images (Ramesh et al., 2021; 2022; Saharia et al., 2022; van den Oord et al., 2016; Salimans et al., 2017; Razavi et al., 2019; Ho et al., 2020; Rombach et al., 2022; Chang et al., 2023; Yu et al., 2024). Recently, scalewise autoregressive models (Tian et al., 2024; Han et al., 2025; Ma et al., 2025) condition from coarse to fine as their progressive sequence, while parallel AR models (Wang et al., 2025; Zhang et al., 2025) select multiple locations at once and condition using sequential blocks.
Both conditioning methods aim to model dependencies between sampling locations. Independently sampling mutually dependent locations can mix modes and produce incompatible values. This is particularly true for adjacent and nearby locations: if an image patch is red, the one next to it is likely to be red as well. Conditioning each location on a common parent, for example in the previous scale, makes the conditional samples more independent of each other (Tian et al., 2024).
Relying exclusively on scalewise conditioning, however, critically depends on a slow scale-up factor: If the scale factor is too large, then an object spanning multiple locations in one scale may not yet be visible in the previous scale, so its appearance dependencies would not be conditioned and independently sampling can mix modes. Rather than scaling up by a factor of 2, as one might consider initially, the best performing multiscale models (Tian et al., 2024; Han et al., 2025; Ma et al., 2025) currently scale by a factor of $\sqrt[3]{2}\approx 1.26$ (i.e. each factor of 2 is further subdivided into 3 evenly spaced scales). However, slow scale-up is not the only way to insert the necessary conditioning. As prior sequential and parallel methods (Wang et al., 2025; Zhang et al., 2025) have shown, another way is to condition between sampling locations.
In this paper, we examine conditioning between locations within each scale, making use of both between- and within-scale conditioning. We develop a sampling order based on a progressive checkerboard, which reduces mutual dependence within each sampling block and enables fast scale-up. Our ordering maintains balance at all levels in a quadtree subdivision, so that varying the block size effectively varies the trade-off between parallelism and modeling conditional dependencies. We use this to explore the relationships between scale-wise and within-scale conditioning, finding somewhat surprisingly that for this spatially balanced setup, the total number of sequential steps largely determines performance independently of how the steps are divided among scales. Our model is competitive with state-of-the-art methods in ImageNet 256x256 class-conditioned generation, using fewer sampling steps.
@@ 2.1 Scalewise Autoregression
Gradual scaling AR models (Tian et al., 2024; Ma et al., 2025; Han et al., 2025), introduced by VAR (Tian et al., 2024), progressively develop a full-resolution array of latent codes, which is used as a working “canvas” to apply residuals. At each iteration, the canvas is downsampled to the current scale (which grows by a factor of $\sqrt[3]{2}$) and a transformer predicts residuals, which are upsampled and applied to the full-resolution array.
@@
Viewing the method this way sheds some light on reasons why the scaling factor remains small. Just as diffusion models struggle with large denoising steps due to their use of independent sampling and Gaussian noise (Song et al., 2021; Salimans and Ho, 2022; Song et al., 2023; Lu et al., 2022), scalewise AR models can struggle with large deblurring steps due to independently sampling underlying multi-modal distributions (Bansal et al., 2023; Hoogeboom and Salimans, 2022).
@@ 2.2 Parallel Sampling
Parallel autoregression (PAR) by Wang et al. (2025) generates images by sampling in parallel from equal-sized square partitions. Their method uses raster order within each partition and no multi-scale sampling, limiting its parallelism to only four groups. Our multi-scale checkerboard sampling uses both multi-scale and within-scale conditioning, as well as an efficient ordering that maintains sample balance and spatial diversity at all levels, reducing the number of sequential steps to just 17, while realizing further performance gains.
More recently, Zhang et al. (2025) use a locality-aware ordering that adds far-away locations while simultaneously growing already-selected regions with adjacent samples, while Pang et al. (2025) use random orderings. However, the former uses a more complex dynamic evolution to grow out regions, while both methods require additional position probe tokens, extending the overall sequence length. We use a simple but effective regular pattern combined with multiscale conditioning, and do not require additional probe tokens.
Yan et al. (2025) use a two-stage "generation then reconstruction" approach applied in the context of MAR (Li et al., 2024) that applies unmasking using diagonal striations, the final stage of which corresponds to a mod 2 checkerboard. In contrast to their method, we use a fully balanced progressive checkerboard, along with explicit multiscale conditioning, in the context of direct autoregression instead of masking. xAR (Ren et al., 2025a) uses flow-matching (Lipman et al., 2023; Liu et al., 2022) to generate large chunks at each AR step. They examine a variety of decompositions, including by scale, finding four quadrants best in their case. Additional recent works include NAR (He et al., 2025), which uses different decoder heads for horizontal and vertical conditioning, leading to a striation pattern, and ARPG (Li et al., 2025), which uses cross-attention for conditioning random-ordered parallel samples with masked models.
Some gradual scaling models also incorporate ways to reduce same-scale dependencies. Han et al. (2025) use random bit flips (“bitwise self-correction”) to address errors from independent sampling at the quantizer bit level after incorporating BSQ (Zhao et al., 2025) as their tokenizer. Ma et al. (2025) use a miniature masked image model applied in the autoregressor head to condition lower-confidence locations on higher-confidence ones. To enable 2x scaling factors, Ren et al. (2025b) stack a flow-matching model on top of mixed-mode AR samples to resolve interdependences. Kutscher et al. (2025) examine the impact of patch ordering in the context of recognition models.
@@ 3.1 Autoregressor
Figure 2: Overview of our multiscale blockwise checkerboard autoregressor.
Our multiscale progressive checkerboard autoregressive model is built around a transformer (Vaswani et al., 2017; Radford et al., 2018; 2019) with blockwise causal mask using successive checkerboard sampling blocks.
@@
During training, the entire sequence is processed in parallel using the ground truth codes $z_{s}$ for both inputs and targets autoregressively. We use a blockwise causal mask where tokens within each block $b_{s}^{i}$ can attend to each other, as well as to tokens from all previous blocks, both for the same scale and in lower-resolution scales. Each token position outputs softmax logits for a quantized codebook; we use flat mean cross-entropy loss over all scales and positions.
Importantly, the locations in each block are spaced as evenly as possible. The number of locations that can be sampled in parallel depends on the independence between locations, when conditioned both on the previous scale and samples in the current scale. Thus, there is a trade-off between the number of scales and the degree of parallelism within each scale, a relation that we explore later in Sec 4.3.
@@ 3.2 Progressive Scan Order
To create the progressive checkerboard ordering, we use a divide-and-conquer approach, where the 2D grid of size $N\times N$ is recursively subdivided into quadrants, and locations selected to maintain a spatially balanced assignment at all quadtree levels. At each recursion step, balanced index lists are generated for each sub-quadrant recursively, then merged using round-robin selection with a diagonal skip-step pattern (i.e., TL, BR, TR, BL). This results in a spatially balanced progressive order, with a unique index in $1\dots N^{2}$ assigned to each of the $N\times N$ spatial positions. (If the grid size $N$ is not a power of 2, we generate the order on the next power of 2 and restrict to the grid size). See Algorithm 1.
This scan order can be used flexibly with various partition sizes. To create the blocks $b_{s}^{i}$ for each scale $s$, we simply divide the progressive scan order into $P$ contiguous segments of equal size. Figure 3 shows an example progressive checkerboard ordering on an $8\times 8$ grid.
@@ 3.3 Token Embeddings
For token embeddings, we use quantized latent codes from a VAE-based autoencoder (Larsen et al., 2016; Esser et al., 2021; Ramesh et al., 2021; Rombach et al., 2022). Each scale is encoded separately from the RGB image, so that latents are directly interpretable and decodable. For transformer inputs, we pass the VAE latent codes through a small MLP to unfold them from their low-dimensional clustering space back to higher dimension. In addition, we use a class token as the first input, which along with AdaLN (Peebles and Xie, 2023) is used to condition on class labels with classifier-free guidance (Ho and Salimans, 2022).
@@ 3.4 Position Encodings and RoPE Mixing
To make use of our blockwise checkerboard order’s ability to model local dependencies, the transformer must be able to attend to adjacent locations for each token. We accomplish this using both input position embeddings and learned rotary encodings (RoPE) (Su et al., 2021; Heo et al., 2024). Embeddings encode the position in the image-scale space, not the sampling order. We initialize RoPE using factored space and scale representations (7/8 spatial, 1/8 scale). The embeddings are then learned in the joint space, allowing the model to find a high-frequency adjacent-position basis and other attention patterns optimized for our sampling order.
Each token in the transformer processes inputs from two locations: the current sampling positions $b_{s}^{i}$, and the conditioning latents at the previous block’s positions $b_{s}^{i-1}$ (Sec. 3.1). Eq. (1) supplies both in the input layer mixture. However, without modification, the attention layers only use RoPEs for the current positions $b_{s}^{i}$. To enable attention on both sets of positions, we looked at mixing RoPEs for attention keys using learned mixing coefficients [^1] We learn the “logits” $\alpha^{\prime}$ of the mixing coefficients and set $\alpha_{lh}={\rm sigmoid}(\alpha^{\prime}_{lh})$. $\alpha_{lh}$ for each transformer layer $l$ and attention head $h$:
@@ 4.1 Model and Training Details
We train our models on ImageNet (Russakovsky et al., 2015) at 256x256 resolution, using two model sizes: the small (S) model has 12 transformer layers with hidden dimension 512 and 16 attention heads, while the large (L) model has 20 layers with 1024 hiddens and 16 heads. In most ablation experiments we use the S model, trained for 100 epochs with batch size 128 and learning rate $1\times 10^{-4}$, dropped to $1\times 10^{-5}$ in two steps over 10 epochs. The L model is trained for 200 epochs with batch size 64 and learning rate $5\times 10^{-5}$, dropped to $1\times 10^{-5}$ for 5 epochs, followed by a second effective drop by increasing batch size to 320 via gradient accumulation for the last 5 epochs. In all cases, we use AdamW (Loshchilov and Hutter, 2019) optimizer with weight decay 0.01 and ten-crop transforms, running on a single NVIDIA GH200 GPU.
We use the VAE autoencoder from LlamaGen (Sun et al., 2024). To better represent smaller image sizes in our multiscale model, we fine-tune by freezing all layers other than the quantizer codebook itself, and retrain the just the codebook layer with a size of 4096 on ImageNet for one epoch, using random image sizes between 16 and 256 and L2 loss on the latent codes.
To measure sample quality, we use Frechet Inception Distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016), computed using 50k samples against the standard reference set using the TensorFlow implementation from Sun et al. (2024); Dhariwal and Nichol (2021).
For classifier-free guidance (CFG), we found applying CFG too early in the sampling progression limits diversity, likely due to the first samples’ influence on global structure. Because of this, we sample the first 5 steps (corresponding to the 1x1 and 2x2 scales when scaling by 2x) with CFG=0, and apply CFG for all subsequent steps. Following common practice (Wang et al., 2025; He et al., 2025; Zhang et al., 2025), we perform a sweep of CFG values at 0.1 increments.
@@ 4.2 Benchmark Comparison
DiT-XL/2 (Peebles and Xie, 2023) Diffu-KL 675M 2.24 278.2 0.83 0.57  $\times 250 11$.9
MAR-L (Li et al., 2024) MAR-KL 479M 1.78 296.0 0.81 0.60 $64\times 100$ 26.4
GtR (Yan et al., 2025) MAR-KL 479M 1.81 297.4 $32\times 30$
xAR (Ren et al., 2025a) Flow-KL 608M 1.28 292.5 0.82 0.62 $4\times 50$ 7.7
LlamaGen-L (Sun et al., 2024) AR-VQ 343M 3.07 256.1 0.83 0.52 576 12.58
VAR-d16 (Tian et al., 2024) AR-VQ 310M 3.30 274.4 0.84 0.51 10 0.12
PAR-L-4x (Wang et al., 2025) AR-VQ 343M 3.76 218.9 0.84 0.50 147 3.38
RandAR-L (Pang et al., 2025) AR-VQ 343M 2.55 288.8 0.81 0.58 88 1.97
NAR-L (He et al., 2025) AR-VQ 372M 3.06 263.9 0.81 0.53 31 1.01
ARPG-L (Li et al., 2025) AR-VQ 320M 2.30 297.7 0.82 0.56 32 0.58
LPD-L (Zhang et al., 2025) AR-VQ 337M 2.40 284.5 0.81 0.57 20 0.28
Checkerboard-L 2x cfg=1.4 (Ours) AR-VQ 343M 2.72 302.5 0.81 0.56 17 0.52
Checkerboard-L 2x cfg=1.5 (Ours) AR-VQ 343M 2.83 318.2 0.82 0.57 17 0.52
Checkerboard-L 4x cfg=1.7 (Ours) AR-VQ 343M 2.79 311.5 0.80 0.57 17 0.52
Table 1: Image generation models on ImageNet 256x256. We mainly compare to other AR-VQ methods (bottom rows) that are directly comparable, but also include non-VQ methods for additional context (top rows, $\times$ indicates mask/ar and diffusion/flow steps). Inference time measured for single image generation on A100 with batch size 1.
@@
Table 1 compares our method against recent autoregression-based image generation models of similar model size on ImageNet 256x256. We show results for our Checkerboard-L model with a 2x scaling ratio and 4 sampling steps per scale (17 total), as well as a 4x ratio with 8 steps per scale (also 17 total) [^2] For the L models, the 2x model is trained from scratch for 200 epochs, and the 4x model is initialized with the 2x model weights and trained for 30 epochs.. Directly relevant comparisons are with other autoregression-based methods with VQ sampling, shown with “AR-VQ” in the second column. We also show methods that use non-quantized autoencoders for wider context, though these are not directly comparable.
Compared to PAR (Wang et al., 2025) and RandAR (Pang et al., 2025), to which our method is most closely related, we achieve similar or better FID and IS with fewer sampling steps and faster inference time. While PAR uses 147 steps and RandAR 88 steps, we only need 17 steps to achieve a FID of 2.72 (compared to 3.76 and 2.55, respectively). Our inference time of 0.52s per image is also faster than both methods (3.38s for PAR and 1.97s for RandAR) on A100. In addition, our method is competitive with the more recent AR-based models ARPG (Li et al., 2025) and LPD (Zhang et al., 2025), notably also using fewer steps.